OpenStudy (anonymous):

anybody know how to solve this? 5sqr x -1=0 i got 1/5 but i have another answer in the book

jimthompson5910 (jim_thompson5910):

5*sqrt(x) - 1 = 0 5*sqrt(x) = 1 sqrt(x) = 1/5 x = (1/5)^2 ... You forgot to square both sides to solve for x x = 1/25

OpenStudy (anonymous):

ok got it thank you :)

jimthompson5910 (jim_thompson5910):

sure thing